OpenStudy (anonymous):

In the function with the equation y =8/x what is the rate of change from x = 1 to x = 2?

OpenStudy (goformit100):

You have to simply differentiate the function and put the respective values.

OpenStudy (anonymous):

The average rate of change of \(x\in (1, 2)\)? That is specified by: \[ \frac{f(b)-f(a)}{b-a} \]Which then becomes: \[ \frac{\frac{8}{2}-\frac{8}{1}}{2-1} \]